# On the reproducibility of discrete-event simulation studies in health research: an empirical study using open models

### v1

@@ 5.1.3 For reporting guidelines
Levels of adherence to the reporting guidelines were similar to those observed in prior studies. The mean proportion of fully met items on the generic reporting checklist42 in this research - either calculated out of applicable criteria (67.4%) or all criteria (61.1%) - was similar to previous evaluations of 211 healthcare DES articles (63.7%)42 and 18 radiotherapy pathway simulation articles (63.6%).93 Gaps in uncertainty assessment, sensitivity analysis, generalisability and validity were common across the studies.42, 93 Although they don’t report on individual items, a previous study did assess adherence to STRESS-DES for 13 pharmacoeconomic models, finding higher mean adherence (80.4%)94 than in this study - as calculated out of applicable criteria (74.0%) or all criteria (68.2%).
In this research, adherence to the reporting checklists did not translate to reproducibility; as observed in prior studies with other checklists.46, 95, 18 This is unsurprising as the generic checklist emphasises model quality whilst STRESS-DES is focused on reporting all necessary information to facilitate replication of simulations. However, some items on these checklists - like describing input parameters, scenarios and dependencies - were helpful in the reproductions. Focusing on STRESS-DES, this could be adapted to further facilitate reproducibility, as has been done with other checklists - for example, the Overview, Design concepts and Details protocol for describing Individual- and Agent-Based Models has been reviewed and updated.96 Suggested improvements for STRESS-DES include:
@@ 5.1.4 For the STARS framework
The STARS framework was designed to support healthcare DES model reuse. As it was published in 2024,43 it was not used in the reviewed studies. This research examined whether incidental adherence to its recommendations improved reproducibility; no clear link was found. This is similar to findings in other fields where practices like documentation and archiving do not guarantee reproducibility.18, 12 This suggests that, while these practices are valuable, other factors contribute to reproducibility challenges. Modifications to STARS that may better facilitate reproducibility include requiring a complete set of materials (i.e. parameters, scenarios, results processing), which aligns with key recommendations from this study and existing badge expectations. Additionally, incorporating guidelines on runtime/hardware, as well as code design (e.g. seeds, running scenarios programmatically, docstrings/comments, output saving).
@@ 5.2 Strengths and limitations
This study is the first to systematically assess the computational reproducibility of healthcare DES models, providing insights that are broadly relevant to other contexts and modelling approaches. A key strength is the transparency of the research, with all code openly licensed, available on GitHub, archived on Zenodo (Appendix C), and supported by a pre-registered protocol.45 Additionally, a complementary website summarises the results and provides further details (https://pythonhealthdatascience.github.io/stars_wp1_summary/).73 The study benefits from a diverse selection of eight studies, although they are biased toward models with shared licensing and resources, which are uncommon in healthcare DES.44 The reproducibility process was time-intensive, requiring up to 28 hours per study, which may not reflect typical researcher resources - but this time investment was crucial to understanding the factors influencing reproducibility, requiring a close examination of the troubleshooting for each case.
Reproductions were conducted by a single researcher, occasionally consulting others when troubleshooting. If the aim of this research had been to determine the exact level of reproducibility, it would have been necessary for multiple researchers to independently attempt to reproduce the studies, to ensure reliable results. A prior study of psychology articles found low inter-rater reliability in such assessments (75% agreement on the executability of R scripts and 56% agreement on reproducibility).21 However, it was felt suitable for a single researcher to conduct the reproductions, given the study’s focus on identifying factors impacting reproducibility, rather than definitive reproduction rates. Barriers and facilitators encountered are expected to be relatively consistent between researchers - as evidenced by the complementary recommendations from reproductions in other fields. Determining successful reproduction involved subjective judgement - acknowledged in the protocol45 - so decisions were made by consensus between at least two team members.
The study focused on models implemented in Python and R, excluding those developed using commercial DES software. Although most DES studies use commercial off-the-shelf software, the majority (62%) of open DES models use FOSS 44. Our expectation is that reproduction of results generated by commercial software may present some differences - particularly in troubleshooting the unique nuances of commercial DES software as well as access to a compatible and suitably licensed version. However, we also expect the top five study recommendations to remain broadly applicable. For example, a simulation model will still need a licence (CC-BY may be best suited to a simulation file); and correct parameters will still be required.
@@ 7 Acknowledgements
We would like to thank the authors who made their code available under open licences, facilitating our research. We are especially grateful to the following authors for their helpful communication and support throughout the project: Names to be added once have received permission from the respective authors.
@@ 8 Data availability statement
The records supporting the findings of this study - including code, models, results and other relevant materials - are synthesised and available at https://pythonhealthdatascience.github.io/stars_wp1_summary/ and archived on Zenodo.73 Links to the individual repositories for the reproduced studies and the original studies are provided in Appendix C.

### v2

@@ 5.1.3 For reporting guidelines
Levels of adherence to the reporting guidelines were similar to those observed in prior studies. The mean proportion of fully met items on the generic reporting checklist41 in this research - either calculated out of applicable criteria (67.4%) or all criteria (61.1%) - was similar to previous evaluations of 211 healthcare DES articles (63.7%)41 and 18 radiotherapy pathway simulation articles (63.6%).91 Gaps in uncertainty assessment, sensitivity analysis, generalisability and validity were common across the studies.41, 91 Although they don’t report on individual items, a previous study did assess adherence to STRESS-DES for 13 pharmacoeconomic models, finding higher mean adherence (80.4%)92 than in this study - as calculated out of applicable criteria (74.0%) or all criteria (68.2%).
In this research, adherence to the reporting checklists did not translate to reproducibility; as observed in prior studies with other checklists.45, 93, 18 This is unsurprising as the generic checklist emphasises model quality whilst STRESS-DES is focused on reporting all necessary information to facilitate replication of simulations. However, some items on these checklists - like describing input parameters, scenarios and dependencies - were helpful in the reproductions. Focusing on STRESS-DES, this could be adapted to further facilitate reproducibility:
@@
Completion and sharing of reporting checklists as supplementary material - though not guaranteed to facilitate reproducibility - certainly increase the transparency of reported models. Alongside the checklists in this research, others exist for different model types - for example, STRESS-ABS for agent-based simulation and STRESS-SD for system dynamics models,2 as well as the ODD protocol (Overview, Design concepts and Details protocol for describing Individual and Agent-Based Models).94 These checklists are complemented by TRACE (TRAnsparent and Comprehensive Ecological modelling documentation), a framework for documenting model development, testing and analysis95 - supported by maintaing modelling notebooks through the development process.96
@@ 5.1.4 For the STARS framework
The STARS framework was designed to support healthcare DES model reuse. As it was published in 2024,42 it was not used in the reviewed studies. This research examined whether incidental adherence to its recommendations improved reproducibility; no clear link was found. This is similar to findings in other fields where practices like documentation and archiving do not guarantee reproducibility.18, 12 This suggests that, while these practices are valuable, other factors contribute to reproducibility challenges. Modifications to STARS that may better facilitate reproducibility include requiring a complete set of materials (i.e. parameters, scenarios, results processing), which aligns with key recommendations from this study and existing badge expectations. Additionally, incorporating guidelines on runtime/hardware, as well as code design (e.g. seeds, running scenarios programmatically, docstrings/comments, output saving).
@@ 5.2 Strengths and limitations
This study is the first to systematically assess the computational reproducibility of healthcare DES models, providing insights that are broadly relevant to other contexts and modelling approaches. A key strength is the transparency of the research, with all code openly licensed, available on GitHub, archived on Zenodo (Appendix C), and supported by a pre-registered protocol.44 Additionally, a complementary website summarises the results and provides further details (https://pythonhealthdatascience.github.io/stars_wp1_summary/).72 The study benefits from a diverse selection of eight studies, although they are biased toward models with shared licensing and resources, which are uncommon in healthcare DES.43 The reproducibility process was time-intensive, requiring up to 28 hours per study, which may not reflect typical researcher resources - but this time investment was crucial to understanding the factors influencing reproducibility, requiring a close examination of the troubleshooting for each case.
Reproductions were conducted by a single researcher, occasionally consulting others when troubleshooting. If the aim of this research had been to determine the exact level of reproducibility, it would have been necessary for multiple researchers to independently attempt to reproduce the studies, to ensure reliable results. A prior study of psychology articles found low inter-rater reliability in such assessments (75% agreement on the executability of R scripts and 56% agreement on reproducibility).21 However, it was felt suitable for a single researcher to conduct the reproductions, given the study’s focus on identifying factors impacting reproducibility, rather than definitive reproduction rates. Barriers and facilitators encountered are expected to be relatively consistent between researchers - as evidenced by the complementary recommendations from reproductions in other fields. Determining successful reproduction involved subjective judgement - acknowledged in the protocol44 - so decisions were made by consensus between at least two team members.
The study focused on models implemented in Python and R, excluding those developed using commercial DES software. Although most DES studies use commercial off-the-shelf software, the majority (62%) of open DES models use FOSS 43. Our expectation is that reproduction of results generated by commercial software may present some differences - particularly in troubleshooting the unique nuances of commercial DES software as well as access to a compatible and suitably licensed version. However, we also expect the top five study recommendations to remain broadly applicable. For example, a simulation model will still need a licence (CC-BY may be best suited to a simulation file); and correct parameters will still be required.
@@ 7 Acknowledgements
We would like to thank the authors who made their code available under open licences, facilitating our research. We are especially grateful to the following authors for their helpful communication during the project: Ivan Hernandez, Tze Ping Loh, Chun Yee Lim, Lois Kim, Kate Johnson, Richard Wood, Mohd Shoaib, and Anastasia Anagnostou. The circular design of Figures 7 and 8 was inspired by Gomes et al. (2022).97
@@ 8 Data availability statement
The records supporting the findings of this study - including code, models, results and other relevant materials - are synthesised and available at https://pythonhealthdatascience.github.io/stars_wp1_summary/ and archived on Zenodo.72 Links to the individual repositories for the reproduced studies and the original studies are provided in Appendix C.
@@ A.4 Journal badges
The criteria for each badge are described in Table A.4, with the badges grouped into three themes, as defined by NISO.62
[…]
@@ Appendix C Code repositories
Links for reproduction and evaluation. Description URL Compilation of results from the eight studies Repository https://github.com/pythonhealthdatascience/stars_wp1_summary Archive 72 Website https://pythonhealthdatascience.github.io/stars_wp1_summary/ Reproduction: Hernandez et al. (2015) 51 Repository https://github.com/pythonhealthdatascience/stars-reproduce-hernandez-2015 Archive 99 Website https://pythonhealthdatascience.github.io/stars-reproduce-hernandez-2015/ Reproduction: Huang et al. (2019) 53 Repository https://github.com/pythonhealthdatascience/stars-reproduce-huang-2019 Archive 100 Website https://pythonhealthdatascience.github.io/stars-reproduce-huang-2019/ Reproduction: Lim et al. (2020) 55 Repository https://github.com/pythonhealthdatascience/stars-reproduce-lim-2020 Archive 101 Website https://pythonhealthdatascience.github.io/stars-reproduce-lim-2020/ Reproduction: Kim et al. (2021) 56 Repository https://github.com/pythonhealthdatascience/stars-reproduce-kim-2021/ Archive 102 Website https://pythonhealthdatascience.github.io/stars-reproduce-kim-2021/ Reproduction: Johnson et al. (2021) 57 Repository https://github.com/pythonhealthdatascience/stars-reproduce-johnson-2021 Archive 103 Website https://pythonhealthdatascience.github.io/stars-reproduce-johnson-2021/ Reproduction: Wood et al. (2021) 50 Repository https://github.com/pythonhealthdatascience/stars-reproduce-wood-2021 Archive 104 Website https://pythonhealthdatascience.github.io/stars-reproduce-wood-2021/ Reproduction: Shoaib et al. (2022) 58 Repository https://github.com/pythonhealthdatascience/stars-reproduce-shoaib-2022 Archive 105 Website https://pythonhealthdatascience.github.io/stars-reproduce-shoaib-2022/ Reproduction: Anagnostou et al. (2022) 60 Repository https://github.com/pythonhealthdatascience/stars-reproduce-anagnostou-2022 Archive 106 Website https://pythonhealthdatascience.github.io/stars-reproduce-anagnostou-2022/
@@
Links to original study repositories. Description URL Original study: Hernandez et al. (2015) 51 Repository https://github.com/ivihernandez/staff-allocation Original study: Huang et al. (2019) 53 Repository https://github.com/shiweih/desECR Original study: Lim et al. (2020) 55 Repository https://github.com/chaose5/COVID-roster-simulation Original study: Kim et al. (2021) 56 Repository https://github.com/mikesweeting/AAA_DES_model Original study: Johnson et al. (2021) 57 Repository https://github.com/KateJohnson/epicR/tree/closed_cohort Original study: Wood et al. (2021) 50 Repository https://github.com/nhs-bnssg-analytics/triage-modelling Original study: Shoaib et al. (2022) 58 Repository https://github.com/shoaibiocl/PHC- Original study: Anagnostou et al. (2022) 60 Repository https://gitlab.com/anabrunel/charm Archive 107